OpenStudy (anonymous):

Solve for x: -2(x + 3) = -2(x + 1) - 4

OpenStudy (anonymous):

i solved it & i am now left with -2x - 6 = -2x - 6 ... is the answer all real numbers or no solution?

OpenStudy (anonymous):

i think you have done something wrong while solving cauze 0=0 is not a solution

OpenStudy (anonymous):

but its okay cus the choices i have for the problem are All real numbers 3 2 No solution

OpenStudy (anonymous):

then you should go with no solution

OpenStudy (anonymous):

you did it right, and go with all real numbers plug a few in and see

OpenStudy (anonymous):

ok im confused, no solution or all real numbers?

OpenStudy (anonymous):

Your equation will finally become f(x)=0 which for every real number has a solution at 0.

OpenStudy (anonymous):

No matter what x you put in, your function will remain 0.

OpenStudy (anonymous):

if \[x \in \mathbb{R}\]

OpenStudy (anonymous):

its no solution cauz its not satisfying left hand side and right hand side.

OpenStudy (anonymous):

It does satisfy left hand and right hand, 0=0 is valid.

OpenStudy (anonymous):

and no matter what real number you put into x, it will satisfy both left and right hand, try it out.

OpenStudy (anonymous):

If you solve the equation it will finally boil down to x=x --> 1=1 --> 0=0

OpenStudy (anonymous):

plug some in and see for yourself... -2(1+3) = -2(1+1)-4 ? -2(97+3) = -2(97+1)-4 ? -2(-5+3) = -2(-5+1)-4 ? try as many as you like, they will all work

OpenStudy (anonymous):

listen bro i mean to say that X is not yielding any value we have to find a particular solution for X we here satisfying the factors for x

OpenStudy (anonymous):

so where does it cross the X axis? Everywhere!!!

OpenStudy (anonymous):

;)

OpenStudy (anonymous):

now i m going to solve your question with every step -2(x+3)=-2(x+1)-4 -2x-6=-2x-2-4 -2x-6=-2x-6 both sides will be cancelled what we have to find was the particular value of x since X has been cancelled from both sides so we got no value of X and so this equation has no solution. And for your info i have asked my teacher too about this problem and teacher said that its a no solution eq